# Is Mamba Effective for Time Series Forecasting?

### v1

@@ Abstract
In the realm of time series forecasting (TSF), the Transformer has consistently demonstrated robust performance due to its ability to focus on the global context and effectively capture long-range dependencies within time, as well as discern correlations between multiple variables. However, due to the inefficiencies of the Transformer model and questions surrounding its ability to capture dependencies, ongoing efforts to refine the Transformer architecture persist. Recently, state space models (SSMs), e.g. Mamba, have gained traction due to their ability to capture complex dependencies in sequences, similar to the Transformer, while maintaining near-linear complexity. In text and image tasks, Mamba-based models can improve performance and cost savings, creating a win-win situation. This has piqued our interest in exploring SSM’s potential in TSF tasks. In this paper, we introduce two straightforward SSM-based models for TSF, S-Mamba and D-Mamba, both employing the Mamba Block to extract variate correlations. Remarkably, S-Mamba and D-Mamba achieve superior performance while saving GPU memory and training time. Furthermore, we conduct extensive experiments to delve deeper into the potential of Mamba compared to the Transformer in the TSF, aiming to explore a new research direction for this field. Our code is available at https://github.com/wzhwzhwzh0921/S-D-Mamba.
@@ Introduction
Time series forecasting (TSF) involves leveraging information from past events to predict conditions in the future. These events often have built-in patterns associated with time or variates, for example, the morning and evening peak patterns in traffic issues, and the pattern between temperature and humidity in weather predictions. This naturally leads to the consideration of identifying long-range dependencies (LRD) or variates correlations (VC) from these data to effectively identify and model these patterns for better forecasting. Transformer (Vaswani et al. 2017), which shines brightly in natural language processing (Wolf et al. 2019, 2020) and visual tasks (Liu et al. 2021; Arnab et al. 2021), is based on the self-attention structure, which can encompass the context and extract deep connections between time points in a sequence. Therefore, it is also applied in TSF tasks. Numerous Transformer-based models with impressive capabilities have been introduced (Wu et al. 2021; Zhou et al. 2022), yet the Transformer architecture faces distinct challenges. Foremost is its quadratic computational complexity, leading to a dramatic increase in calculations for long input sequences. Numerous models have attempted to reduce computational complexity by modifying the Transformer’s structure, such as focusing only on a portion of the sequence (Kitaev, Kaiser, and Levskaya 2020; Zhou et al. 2021; Li et al. 2019). However, the loss of node information in this process may also lead to certain performance degradations. A more promising approach involves using linear models (Li et al. 2023a; Zeng et al. 2023), which possesses linear computational complexity. However, compared with the state-of-the-art Transformer models, these models do not perform so well due to their inability to capture LRD or VC. Additionally, since they do not predict trends and rely solely on numerical calculations, accurate predictions can only be achieved when sufficient input information is available (Zeng et al. 2023).
@@
It seems that capturing browsing context, building LRDs or VCs and linear complexity is not achievable simultaneously. The State Space Models (SSM) (Gu, Goel, and Ré 2021; Smith, Warrington, and Linderman 2022) demonstrate potential in simultaneously optimizing both aspects. It employs convolutional calculation to capture LRDs and eliminates hidden states making it benefit from parallel computing and achieving approximately linear complexity in processing speed. Consequently, SSM is inherently well-suited for TSF tasks. A Previous study (Rangapuram et al. 2018) also attempted to employ SSM for TSF, but the architecture is unable to identify and filter content effectively, and the dependencies it captures are solely based on distance, resulting in suboptimal performance. Recent work, Mamba (Gu and Dao 2023), introduces a selective mechanism into SSM, enabling it to discern the importance of information like the attention mechanism. This development has prompted researchers to build Mamba-based models in domains where Transformer has demonstrated superior performance. The outcomes indicate that Mamba exhibits significant potential in both text and image fields, frequently achieving a win-win situation in terms of model performance and computational efficiency (Zhu et al. 2024; Yang, Xing, and Zhu 2024). This improvement is attributed to Mamba’s enhanced capability to process long sequence data, making it inherently suitable for TSF tasks and sparking our interest in exploring Mamba’s application in this area.
As a result, we launch Simple-Mamba (S-Mamba) and Dual-Mamba (D-Mamba) in an attempt to reconcile the performance and speed disputes in the TSF field and open up new research directions for it. S-Mamba employs one Mamba block to process VC, while D-Mamba incorporates an additional mamba block compared to S-Mamba for VC. The distinction between the additional Mamba block in D-Mamba and the original Mamba block lies in the former’s sensitivity to variable correlations (VC) is higher by adjusting the parameters of the Mamba block. In our design, this modification serves two purposes. Firstly, it is akin to multi-head attention mechanism, aiming to capture multiple facets of VC. Secondly, it is intended to enable the model to learn higher-level VC based on the two different sensitivities. Ultimately, our experimental results demonstrate that S-Mamba and D-Mamba not only achieve significant savings in GPU memory usage and training time but also maintain superior performance compared to state-of-the-art models. our contributions can be categorized into the following three aspects:
  • We introduce S-Mamba and D-Mamba, utilizing the SSM mechanism to capture variate correlations. S-Mamba only has one Mamba block, while D-Mamba uses two Mamba blocks with different sensitivity.
@@
  • We evaluate of S-Mamba and D-Mamba. They not only reduce the time and GPU memory required for forecasts but also maintain superior performance.
@@
  • We conduct extensive experiments to further delve deeper into Mamba’s potential in TSF tasks, aiming to explore a new research direction in this field.
@@
However, to the best of our knowledge, there is still no work applying the Mamba architecture to time series forecasting tasks. Intuitively, the Mamba architecture effectively mitigates the problems faced by the Transformer and MLP architectures for time series forecasting. Therefore, we try to introduce the Mamba architecture in time series forecasting tasks.
@@
Mamba (Gu and Dao 2023) introduces a data-dependent selection mechanism into the S4 and incorporates hardware-aware parallel algorithms in its looping mode, enabling Mamba to effectively capture contextual information in long sequences while maintaining computational efficiency. As an approximately linear perplexity series model, Mamba demonstrates significant potential in long sequence tasks, compared to transformers, in both efficiency enhancement and performance improvement.
Figure 1: Overall framework of S/D-Mamba, the left side of the figure presents the overall architecture of our model. The Mamba (sensitive) enclosed within the dotted line represents an additional Mamba block included in D-Mamba. The right side of the figure details the components of the S/D-Mamba Block.
@@
In this section, we provide a detailed introduction for S/D-Mamba. Fig.1 illustrates the overall structure of S/D-Mamba, which is primarily composed of three layers. The first layer, the embedding layer, utilizes a variable-based embedding approach. The second layer, the Mamba layer (S/D-Mamba block), forms the core of our model, capturing the dependencies among input variates’ embeddings and updating them according to variate correlations. The third layer, the Norm-FFN-Norm layer, further mixes the information output by the Mamba layer. Alg.1 demonstrates the operation process of S-Mamba. The procedure of D-Mamba remains unchanged, except for the addition of a parallel sensitive Mamba block within the Mamba Layer.
@@
1:  Embedding Layer:
2:  $Batch(U_{in}^{\top}):(B,V,L)\leftarrow\mathrm{Transpose}(Batch(U_{in}))$Turn to variate side}
3:  $\bm{U}:(B,V,D)\leftarrow\mathrm{Linear}(Batch(U_{in}^{\top}))$Variate-based Embedding}
4:  for $l$ in $Layers$ do
5:     S-Mamba Layer:
6:     $x,z:(B,V,ED)\leftarrow\mathrm{Linear}(\bm{U})$Linear projection}
7:     $x^{{}^{\prime}}:(B,V,ED)\leftarrow\mathrm{SiLU}(\mathrm{Conv1D}(x))$
8:     $\textbf{\emph{A}}:(D,N)\leftarrow Parameter$Structured state matrix}
[…]
10:     $\Delta:(B,V,D)\leftarrow\mathrm{Softplus}(Parameter+\mathrm{Broadcast}(\mathrm{Linear}(x^{{}^{\prime}})))$
11:     $:(B.V.D.N)\leftarrow{discretize}(\Delta,\textbf{\emph{A}},\textbf{\emph{B}})$ A , B ) {Input-dependent parameters and discretization}
12:     $y:(B,V,ED)\leftarrow\mathrm{SelectiveSSM}(\overline{\textbf{\emph{A}}},\overline{\textbf{\emph{B}}},\textbf{\emph{C}})(x^{{}^{\prime}})$
13:     $y^{{}^{\prime}}:(B,V,ED)\leftarrow y\otimes\mathrm{SiLU}(z)$
14:     $\bm{Y}:(B,V,D)\leftarrow\mathrm{Linear}(y^{{}^{\prime}})$Linear Projection}
15:     Norm-FFN-Norm Layer:
16:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{Dropout}(\bm{Y})+\bm{U}$Residual Connection}
17:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{LayerNorm}(\bm{U}^{{}^{\prime}})$
18:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{Feed-forward}(\bm{U}^{{}^{\prime}})$
19:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{LayerNorm}(\bm{U}^{{}^{\prime}})$
20:  end for
21:  Projection:
22:  $\bm{U}^{{}^{\prime}}:(B,V,T)\leftarrow\mathrm{Projection}(\bm{U}^{{}^{\prime}})$
23:  $Batch(U_{out}):(B,T,V)\leftarrow\mathrm{Transpose}(\bm{U}^{{}^{\prime}})$
Embedding Layer
The input for the Embedding layer is $U_{in}$. Similar to iTransformer (Liu et al. 2023), we also set up a variate-based linear embedding layer, which considers each variate independently, thereby effectively reducing the noise caused by mixed variates according to time when embedding. After that, each variate information of the input data point is mapped into a high-dimensional space, where $D$ denotes the embedding dimension as (3).
@@
S/D-Mamba layer
The Mamba layer is considered the fundamental building block of the model architecture proposed in this paper. It functions to fuse information between variables based on VC.
Figure 2: The structure of selective SSM.
Mamba Block. The details are presented in the algorithm related to the mamba layer in Alg.1 and the description in Fig. 2. The Mamba layer takes a sequence $\bm{U}\in\mathbb{R}^{B\times V\times D}$ as input at time step $t$, where $B$ denotes the batch size, $V$ denotes the number of variates, and $D$ denotes hidden dimension. The block first expands the hidden dimension to $ED$ through linear projection, obtaining $x$ and $z$. Then, it processes the projection obtained earlier using convolutional functions and a SiLU (Elfwing, Uchibe, and Doya 2017) activation function. Based on the discretized SSM selected by the input parameters, denoted as the core of the mamba block, together with $x^{\prime}$, it generates the state representation $y$. Finally, $y$ is combined with a residual connection from $z$ after activation, and the final output $y_{t}$ at time step $t$ is obtained through a linear transformation. In summary, the Mamba Block effectively handles sequential information by leveraging selective state space models and input-dependent adaptations. The parameters involved in the Mamba Block include an SSM state expansion factor $N$, a size of convolutional kernel $k$, and a block expansion factor $E$ for input-output linear projection. The larger the values of $N$ and $E$, the higher the computational cost. The final output of the Mamba block is $\bm{Y}\in\mathbb{R}^{B\times V\times D}$. When there are two Mamba Blocks (D-Mamba), the inputs of the two mamba blocks are aggregated as the output $\bm{Y}=\bm{Y_{1}}+\bm{Y_{2}}$, where $\bm{Y_{1}}$ and $\bm{Y_{2}}$ are the outputs of the two mamba layers respectively.
Norm-FFN-Norm Layer
@@
At this layer, we further process the output of the Mamba Layer. Firstly, we employ a normalization layer (Liu et al. 2023), originally designed to enhance convergence and training stability in deep networks. Prior to layer normalization, we standardize all variables to a Gaussian distribution, thereby minimizing disparities resulting from inconsistent measurements. Then, the feed-forward network (FFN) is used on the series representation of each variate, allowing for complex time series descriptions. FFN encodes observed time series and decodes future series representations using dense non-linear connections, similar to recent MLP-based approaches. Finally, another normalization layer same as the former is set to adjust the future series representations.
Table 1: Full results of the long-term time series forecasting task. We compare the representative and the state-of-the-art models with S-Mamba and D-Mamba under different prediction lengths following the setting of iTransformer. The lookback length $L$ is set to 96 for all models. The number on the left side of the table represents the length of the output sequence $T$. The optimal results are highlighted in bold red font, while the suboptimal results are presented in underlined blue font. The results of baselines are reported by Liu et al. (2023).
Models D-Mamba S-Mamba iTransformer RLinear PatchTST DLinear FEDformer Autoformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03 12 0.066 0.171 0.066 0.171 0.071 0.174 0.126 0.236 0.099 0.216 0.122 0.243 0.126 0.251 0.272 0.385
24 0.087 0.196 0.088 0.197 0.093 0.201 0.246 0.334 0.142 0.259 0.201 0.317 0.149 0.275 0.334 0.440
48 0.133 0.243 0.138 0.248 0.125 0.236 0.551 0.529 0.211 0.319 0.333 0.425 0.227 0.348 1.032 0.782
96 0.197 0.301 0.217 0.322 0.164 0.275 1.057 0.787 0.269 0.370 0.457 0.515 0.348 0.434 1.031 0.796
PEMS04 12 0.073 0.178 0.075 0.179 0.078 0.183 0.138 0.252 0.105 0.224 0.148 0.272 0.138 0.262 0.424 0.491
24 0.085 0.193 0.087 0.196 0.095 0.205 0.258 0.348 0.153 0.275 0.224 0.340 0.177 0.293 0.459 0.509
48 0.104 0.214 0.105 0.217 0.120 0.233 0.572 0.544 0.229 0.339 0.355 0.437 0.270 0.368 0.646 0.610
96 0.125 0.234 0.128 0.239 0.150 0.262 1.137 0.820 0.291 0.389 0.452 0.504 0.341 0.427 0.912 0.748
PEMS07 12 0.062 0.160 0.064 0.162 0.067 0.165 0.118 0.235 0.095 0.207 0.115 0.242 0.109 0.225 0.199 0.336
24 0.079 0.183 0.081 0.184 0.088 0.190 0.242 0.341 0.150 0.262 0.210 0.329 0.125 0.244 0.323 0.420
48 0.091 0.188 0.094 0.193 0.110 0.215 0.562 0.541 0.253 0.340 0.398 0.458 0.165 0.288 0.390 0.470
96 0.197 0.301 0.199 0.305 0.139 0.245 1.096 0.795 0.346 0.404 0.594 0.553 0.262 0.376 0.554 0.578
PEMS08 12 0.084 0.189 0.085 0.189 0.079 0.182 0.133 0.247 0.168 0.232 0.154 0.276 0.173 0.273 0.436 0.485
24 0.126 0.232 0.127 0.233 0.115 0.219 0.249 0.343 0.224 0.281 0.248 0.353 0.210 0.301 0.467 0.502
48 0.180 0.228 0.196 0.229 0.186 0.235 0.569 0.544 0.321 0.354 0.440 0.470 0.320 0.394 0.966 0.733
96 0.243 0.254 0.245 0.255 0.221 0.267 1.166 0.814 0.408 0.417 0.674 0.565 0.442 0.465 1.385 0.915
ETTm1 96 0.329 0.364 0.325 0.361 0.334 0.368 0.355 0.376 0.329 0.367 0.345 0.372 0.379 0.419 0.505 0.475
192 0.371 0.387 0.368 0.385 0.377 0.391 0.391 0.392 0.367 0.385 0.380 0.389 0.426 0.441 0.553 0.496
336 0.403 0.409 0.401 0.408 0.426 0.420 0.424 0.415 0.399 0.410 0.413 0.413 0.445 0.459 0.621 0.537
720 0.471 0.448 0.469 0.446 0.491 0.459 0.487 0.450 0.454 0.439 0.474 0.453 0.543 0.490 0.671 0.561
ETTm2 96 0.179 0.262 0.180 0.264 0.180 0.264 0.182 0.265 0.175 0.259 0.193 0.292 0.203 0.287 0.255 0.339
192 0.245 0.307 0.245 0.306 0.250 0.309 0.246 0.304 0.241 0.302 0.284 0.362 0.269 0.328 0.281 0.340
336 0.309 0.347 0.309 0.347 0.311 0.348 0.307 0.342 0.305 0.343 0.369 0.427 0.325 0.366 0.339 0.372
720 0.406 0.402 0.408 0.403 0.412 0.407 0.407 0.398 0.402 0.400 0.554 0.522 0.421 0.415 0.433 0.432
ETTh1 96 0.385 0.405 0.382 0.401 0.386 0.405 0.386 0.395 0.414 0.419 0.386 0.400 0.376 0.419 0.449 0.459
192 0.433 0.433 0.434 0.435 0.441 0.436 0.437 0.424 0.460 0.445 0.437 0.432 0.420 0.448 0.500 0.482
336 0.475 0.455 0.482 0.458 0.487 0.458 0.479 0.446 0.501 0.466 0.481 0.459 0.459 0.465 0.521 0.496
720 0.482 0.479 0.481 0.480 0.503 0.491 0.481 0.470 0.500 0.488 0.519 0.516 0.506 0.507 0.514 0.512
ETTh2 96 0.294 0.348 0.295 0.348 0.297 0.349 0.288 0.338 0.302 0.348 0.333 0.387 0.358 0.397 0.346 0.388
192 0.376 0.398 0.376 0.397 0.380 0.400 0.374 0.390 0.388 0.400 0.477 0.476 0.429 0.439 0.456 0.452
336 0.417 0.430 0.420 0.432 0.428 0.432 0.415 0.426 0.426 0.433 0.594 0.541 0.496 0.487 0.482 0.486
720 0.434 0.449 0.425 0.441 0.427 0.445 0.420 0.440 0.431 0.446 0.831 0.657 0.463 0.474 0.515 0.511
Electricity 96 0.142 0.238 0.144 0.239 0.148 0.240 0.201 0.281 0.181 0.270 0.197 0.282 0.193 0.308 0.201 0.317
192 0.160 0.253 0.160 0.252 0.162 0.253 0.201 0.283 0.188 0.274 0.196 0.285 0.201 0.315 0.222 0.334
336 0.176 0.271 0.176 0.271 0.178 0.269 0.215 0.298 0.204 0.293 0.209 0.301 0.214 0.329 0.231 0.338
720 0.203 0.297 0.207 0.300 0.225 0.317 0.257 0.331 0.246 0.324 0.245 0.333 0.246 0.355 0.254 0.361
Exchange 96 0.088 0.209 0.087 0.209 0.086 0.206 0.093 0.217 0.088 0.205 0.088 0.218 0.148 0.278 0.197 0.323
192 0.184 0.306 0.182 0.304 0.177 0.299 0.184 0.307 0.176 0.299 0.176 0.315 0.271 0.315 0.300 0.369
336 0.334 0.419 0.336 0.421 0.331 0.417 0.351 0.432 0.301 0.397 0.313 0.427 0.460 0.427 0.509 0.524
720 0.866 0.702 0.857 0.698 0.847 0.691 0.886 0.714 0.901 0.714 0.839 0.695 1.195 0.695 1.447 0.941
Traffic 96 0.387 0.260 0.387 0.262 0.395 0.268 0.649 0.389 0.462 0.295 0.650 0.396 0.587 0.366 0.613 0.388
192 0.408 0.269 0.410 0.270 0.417 0.276 0.601 0.366 0.466 0.296 0.598 0.370 0.604 0.373 0.616 0.382
336 0.423 0.276 0.425 0.277 0.433 0.283 0.609 0.369 0.482 0.304 0.605 0.373 0.621 0.383 0.622 0.337
720 0.458 0.294 0.459 0.296 0.467 0.302 0.647 0.387 0.514 0.322 0.645 0.394 0.626 0.382 0.660 0.408
Weather 96 0.170 0.212 0.172 0.214 0.174 0.214 0.192 0.232 0.177 0.218 0.196 0.255 0.217 0.296 0.266 0.336
192 0.220 0.257 0.221 0.259 0.221 0.254 0.240 0.271 0.225 0.259 0.237 0.296 0.276 0.336 0.307 0.367
336 0.276 0.298 0.280 0.300 0.278 0.296 0.292 0.307 0.278 0.297 0.283 0.335 0.339 0.380 0.359 0.395
720 0.356 0.348 0.358 0.349 0.358 0.347 0.364 0.353 0.354 0.348 0.345 0.381 0.403 0.428 0.419 0.428
Solar 96 0.202 0.233 0.208 0.241 0.203 0.237 0.322 0.339 0.234 0.286 0.290 0.378 0.242 0.342 0.884 0.711
192 0.246 0.263 0.232 0.259 0.233 0.261 0.359 0.356 0.267 0.310 0.320 0.398 0.285 0.380 0.834 0.692
336 0.257 0.263 0.262 0.281 0.248 0.273 0.397 0.369 0.290 0.315 0.353 0.415 0.282 0.376 0.941 0.723
720 0.256 0.283 0.258 0.284 0.249 0.275 0.397 0.356 0.289 0.317 0.356 0.413 0.357 0.427 0.882 0.717
$1^{\text{st}}$ Count 22 19 7 13 9 6 5 10 9 7 3 0 3 0 0 0
$2^{\text{nd}}$ Count 18 19 23 23 9 8 1 1 3 5 1 3 0 1 0 0

### v2

@@ Abstract
In the realm of time series forecasting (TSF), it is imperative for models to adeptly discern and distill dependencies embedded within historical time series data. This encompasses the extraction of temporal dependencies and inter-variate correlations (VC), thereby empowering the models to forecast future states. Transformer-based models have exhibited formidable efficacy in TSF, primarily attributed to their distinct proficiency in apprehending both TD and VC. However, due to the inefficiencies, ongoing efforts to refine the Transformer persist. Recently, state space models (SSMs), e.g. Mamba, have gained traction due to their ability to process complex dependencies in sequences, similar to the Transformer, while maintaining near-linear complexity. This has piqued our interest in exploring SSM’s potential in TSF tasks. Therefore, we propose a Mamba-based model named Simple-Mamba (S-Mamba) for TSF. Specifically, we tokenize the time points of each variate autonomously via a linear layer. Subsequently, a bidirectional Mamba layer is utilized to extract VC, followed by the generation of forecast outcomes through a composite structure of a Feed-Forward Network for TD and a mapping layer. Experiments on several datasets prove that S-Mamba maintains low computational overhead and achieves leading performance. Furthermore, we conduct extensive experiments to delve deeper into the potential of Mamba compared to the Transformer in the TSF. Our code is available at https://github.com/wzhwzhwzh0921/S-D-Mamba.
@@ Introduction
Time series forecasting (TSF) involves leveraging information from past events to predict conditions in the future. These events often have built-in patterns associated with time or variates, for example, the morning and evening peak patterns in traffic issues, and the pattern between temperature and humidity in weather predictions. This naturally leads to the consideration of identifying temporal dependencies (TD) and inter-variate correlations (VC) from these data to effectively model these patterns for better forecasting.
@@
Transformer (Vaswani et al. 2017), which shines brightly in natural language processing (Wolf et al. 2019, 2020) and visual tasks (Liu et al. 2021; Arnab et al. 2021), is based on the self-attention structure, which can encompass the context and extract deep connections between time points in a sequence. Therefore, it is also applied in TSF tasks. Numerous Transformer-based models with impressive capabilities have been introduced (Wu et al. 2021; Zhou et al. 2022), yet the Transformer architecture faces distinct challenges. Foremost is its quadratic computational complexity, leading to a dramatic increase in calculations for long input sequences. Numerous models have attempted to reduce computational complexity by modifying the Transformer’s structure, such as focusing only on a portion of the sequence (Kitaev, Kaiser, and Levskaya 2020; Zhou et al. 2021; Li et al. 2019). However, the loss of information in this process may also lead to certain performance degradations. A more promising approach involves using linear models (Li et al. 2023a; Zeng et al. 2023), which possesses linear computational complexity. However, since they do not learn in-context and rely solely on linear numerical calculations, their performance is suboptimal compared to state-of-the-art Transformer models and accurate forecasts can only be achieved when sufficient input information is available (Zeng et al. 2023).
@@
The State Space Models (SSM) (Gu, Goel, and Ré 2021; Smith, Warrington, and Linderman 2022) demonstrate potential in simultaneously optimizing performance and computational complexity. It employs convolutional calculation to capture sequence information and eliminates hidden states making it benefit from parallel computing and achieving near-linear complexity in processing speed. Consequently, SSM is inherently well-suited for TSF tasks. A Previous study (Rangapuram et al. 2018) also attempted to employ SSM for TSF, but the SSM architecture it used is unable to identify and filter content effectively, and the dependencies it captures are solely based on distance, resulting in unsatisfactory performance. Recent work, Mamba (Gu and Dao 2023), introduces a selective mechanism into SSM, enabling it to discern the importance of information like the attention mechanism. This development has prompted researchers to build Mamba-based models in domains where Transformer has demonstrated superior performance. The outcomes indicate that Mamba exhibits significant potential in both text and image fields, frequently achieving a win-win situation in terms of model performance and computational efficiency (Zhu et al. 2024; Yang, Xing, and Zhu 2024). This improvement is attributed to Mamba’s enhanced capability to process long sequence data, making it inherently suitable for TSF tasks and sparking our interest in exploring Mamba’s application in this field.
As a result, we launch a Mamba-based model Simple-Mamba (S-Mamba) in an attempt to reconcile the performance and speed disputes for TSF. In the S-Mamba, the time points of each variate are tokenized by a linear layer. Subsequently, a Mamba VC Encoding layer is employed to encode the VC to leverage the inter-variate mutual information for augmenting the forecasting. A FFN TD Encoding Layer containing a simple Feed-Forward Network (FFN) is followed to extract the TD. Ultimately, a mapping layer is utilized to output the forecast results. Experimental results demonstrate that S-Mamba not only achieves savings in GPU memory usage and training time but also maintains superior performance compared to state-of-the-art models. Concurrently, extensive experiments are conducted to evaluate Mamba’s efficacy and potential in TSF, for instance, to evaluate whether Mamba possesses equivalent generalization capabilities as the Transformer. Our contributions can be categorized into the following three aspects:
  • We propose S-Mamba, a Mamba-based model for TSF, which delegates the extraction of inter-variate correlations and temporal dependencies to a bidirectional Mamba block and a Feed-Forward network.
@@
  • We evaluate the performance of S-Mamba, which not only has low GPU memory required and short time for forecasts but also maintains superior performance compared to the representative and state-of-the-art models.
@@
  • We conduct extensive experiments to further delve deeper into Mamba’s potential in TSF tasks.
@@
Intuitively, the Mamba architecture can effectively mitigate the problems faced by the Transformer and MLP architectures for time series forecasting. Therefore, we try to introduce the Mamba architecture in time series forecasting tasks.
@@
Algorithm 1 The process of Mamba Block
Input: $\bm{X}:(B,V,D)$
Output: $\bm{Y}:(B,V,D)$
1:  $x,z:(B,V,ED)\leftarrow\mathrm{Linear}(\bm{U})$Linear projection}
2:  $x^{{}^{\prime}}:(B,V,ED)\leftarrow\mathrm{SiLU}(\mathrm{Conv1D}(x))$
3:  $\textbf{\emph{A}}:(D,N)\leftarrow Parameter$Structured state matrix}
[…]
5:  $\Delta:(B,V,D)\leftarrow\mathrm{Softplus}(Parameter+\mathrm{Broadcast}(\mathrm{Linear}(x^{{}^{\prime}})))$
6:  $:(B.V.D.N)\leftarrow{discretize}(\Delta,\textbf{\emph{A}},\textbf{\emph{B}})$ A , B ) {Input-dependent parameters and discretization}
7:  $y:(B,V,ED)\leftarrow\mathrm{SelectiveSSM}(\overline{\textbf{\emph{A}}},\overline{\textbf{\emph{B}}},\textbf{\emph{C}})(x^{{}^{\prime}})$
8:  $y^{{}^{\prime}}:(B,V,ED)\leftarrow y\otimes\mathrm{SiLU}(z)$
9:  $\bm{Y}:(B,V,D)\leftarrow\mathrm{Linear}(y^{{}^{\prime}})$Linear Projection}
Figure 1: The structure of selective SSM (Mamba).
Mamba Block
Mamba (Gu and Dao 2023) introduces a data-dependent selection mechanism into the S4 and incorporates hardware-aware parallel algorithms in its looping mode, enabling Mamba to effectively capture contextual information in long sequences while maintaining computational efficiency. As an approximately linear perplexity series model, Mamba demonstrates significant potential in long sequence tasks, compared to transformers, in both efficiency enhancement and performance improvement. The details are presented in the algorithm related to the mamba layer in Alg.1 and the description in Fig. 1, where the former illustrates the complete data processing procedure, while the latter depicts the formation process of the output at sequence position $t$. The Mamba layer takes a sequence $\bm{X}\in\mathbb{R}^{B\times V\times D}$ as input, where $B$ denotes the batch size, $V$ denotes the number of variates, and $D$ denotes hidden dimension. The block first expands the hidden dimension to $ED$ through linear projection, obtaining $x$ and $z$. Then, it processes the projection obtained earlier using convolutional functions and a SiLU (Elfwing, Uchibe, and Doya 2017) activation function to get $x^{\prime}$. Based on the discretized SSM selected by the input parameters, denoted as the core of the mamba block, together with $x^{\prime}$, it generates the state representation $y$. Finally, $y$ is combined with a residual connection from $z$ after activation, and the final output $y_{t}$ at time step $t$ is obtained through a linear transformation. In summary, the Mamba Block effectively handles sequential information by leveraging selective state space models and input-dependent adaptations. The parameters involved in the Mamba Block include an SSM state expansion factor $N$, a size of convolutional kernel $k$, and a block expansion factor $E$ for input-output linear projection. The larger the values of $N$ and $E$, the higher the computational cost. The final output of the Mamba block is $\bm{Y}\in\mathbb{R}^{B\times V\times D}$.
Figure 2: Overall framework of S-Mamba, the left side of the figure presents the overall architecture of our model. The right side of the figure details the components of the S-Mamba Block.
@@
In this section, we provide a detailed introduction of S-Mamba. Fig.2 illustrates the overall structure of S-Mamba, which is primarily composed of four layers. The first layer, the Linear Tokenization Layer, tokenizes the time series with a linear layer. The second layer, the Mamba inter-Variate correlation (VC) Encoding layer, employs a bidirectional Mamba block to capture mutual information among variates. The third layer, the FFN Temporal Dependencies (TD) Encoding Layer, further learns the temporal sequence information and finally generates future series representations by a Feed-Forward Network. Then the final layer, the Projection Layer, is only responsible for mapping the processed information of the above layers as the model’s forecast. Alg.2 demonstrates the operation process of S-Mamba.
@@
1:  Linear Tokenization Layer:
2:  $Batch(U_{in}^{\top}):(B,V,L)\leftarrow\mathrm{Transpose}(Batch(U_{in}))$
3:  $\bm{U}^{tok}:(B,V,D)\leftarrow\mathrm{LinearTokenize}(Batch(U_{in}^{\top}))$Tokenization}
4:  for $l$ in $Mamba~{}Layers$ do
5:     Mamba VC Encoding Layer:
6:     $\overrightarrow{\bm{Y}}:(B,V,D)\leftarrow\overrightarrow{\mathrm{Mamba~{}Block}}(\bm{U})$
7:     $\overleftarrow{\bm{Y}}:(B,V,D)\leftarrow\overleftarrow{\mathrm{Mamba~{}Block}}(\bm{U})$
8:     $\bm{Y}:(B,V,D)\leftarrow\overrightarrow{\bm{Y}}+\overleftarrow{\bm{Y}}$Fusion Bidirectional Information}
9:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\bm{Y}+\bm{U}$Residual Connection}
10:     FFN TD Encoding Layer:
11:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{LayerNorm}(\bm{U}^{{}^{\prime}})$
12:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{Feed-Forward}(\bm{U}^{{}^{\prime}})$
13:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{LayerNorm}(\bm{U}^{{}^{\prime}})$
14:  end for
15:  Projection:
16:  $\bm{U}^{{}^{\prime}}:(B,V,T)\leftarrow\mathrm{Projection}(\bm{U}^{{}^{\prime}})$
17:  $Batch(U_{out}):(B,T,V)\leftarrow\mathrm{Transpose}(\bm{U}^{{}^{\prime}})$
Linear Tokenization Layer
The input for the Embedding layer is $U_{in}$. Similar to iTransformer (Liu et al. 2023), we commence by tokenizing the time series, a method analogous to the tokenization of sequential text in natural language processing, to standardize the temporal series format. This pivotal task is executed by a single linear layer (3).
@@
where $\bm{U}$ is the output of this layer.
Mamba VC Encoding Layer
Within this layer, our primary objective is to extract the VC by linking variates that exhibit analogous trends aiming to learn the mutual information therein, thereby enabling a more robust representation of variates. The Transformer architecture confers the capacity for global attention (Vaswani et al. 2017), enabling the computation of the impact of all other variables upon a given variable, which facilitates the learning of precise information. Nonetheless, with an escalation in the number of variables, the computational burden of this global attention—characterized by its quadratic complexity—becomes untenable. In contrast, Mamba’s selective mechanism can discern the significance of different variables akin to an attention mechanism, and it exhibits a computational overhead that escalates in a near-linear fashion with an increasing count of variables. Yet, the unilateral nature of Mamba precludes it from attending to global variables in the manner of the Transformer; its selection mechanism is unidirectional, capable only of incorporating antecedent variables. To surmount this limitation, we employ two Mamba blocks to be combined as a bidirectional Mamba layer (4), which facilitates the acquisition of correlations among all variables, both antecedent and subsequent.
$\overrightarrow{\bm{Y}}=\overrightarrow{\mathrm{Mamba~{}Block}}(\bm{U})$ (4)
$\overleftarrow{\bm{Y}}=\overleftarrow{\mathrm{Mamba~{}Block}}(\bm{U})$
The VC encoded by the bidirectional Mamba is aggregated $\bm{Y}=\overrightarrow{\bm{Y}}+\overleftarrow{\bm{Y}}$ and connected with another residual network to form the output of this layer $\bm{U}^{{}^{\prime}}=\mathrm{\bm{Y}}+\bm{U}$.
FFN TD Encoding Layer
@@
At this layer, we further process the output of the Mamba VC Encoding Layer. Firstly, we employ a normalization layer (Liu et al. 2023) to enhance convergence and training stability in deep networks by standardizing all variates to a Gaussian distribution, thereby minimizing disparities resulting from inconsistent measurements. Then, the feed-forward network (FFN) is used on the series representation of each variate. The FFN layer encodes observed time series and decodes future series representations using dense non-linear connections. During this procedure, FFN implicitly encodes TD. Finally, another normalization layer is set to adjust the future series representations as line 11-13 in Alg.1.
Projection Layer
Based on the output of the FFN TD Encoding layer, the tokenized temporal information is reconstructed into the time series requiring prediction via a mapping layer, subsequently undergoing transposition to yield the final predictive outcome.